OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. A. -three fourthsx + y = 10 B. 3x - 4y = -40 C. y = three fourthsx + 12 D. y = three fourthsx + 10

OpenStudy (anonymous):

\[y-4=\frac{ 3 }{ 4 }\left( x+8 \right)\] 4y-16=3x+24 simplify it.

OpenStudy (anonymous):

B